OpenStudy (anonymous):

y=5x+7 in a form of Ax+By=C

OpenStudy (ahsome):

You want the \(x\) and \(y\) value in one section, and the rest in another. Easy way of doing that is to move the \(5x\) from the Right Hand Side to the Left Hand Side \[y=5x+7\]\[y-4x=7\] Then, times everything by \(-1\) to get the plus in the middle \[y-4x=7\]\[-1(y-4x)=-1(7)\]\[-y+4x=-7\]

OpenStudy (ahsome):

Does that make sense, @Cutiecutie?

OpenStudy (ahsome):

If you think I helped @Cutiecutie , please press the Best Response button. Thanks :D

OpenStudy (anonymous):

"The standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers." 5x - y = -7